Maslyn Manufacturing has projected sales of $148.2 million next year. Costs are expected to be $82.6 million, and net investment is expected to be $16.6 million. Each of these values is expected to grow at 14 percent the following year, with the growth rate declining by 2 percent per year until the growth rate reaches 6 percent, where it is expected to remain indefinitely. There are 7.1 million shares of stock outstanding and investors require a return of 13 percent on the company’s stock. The corporate tax rate is 39 percent.

a. What is your estimate of the current stock price?

b. Suppose instead that you estimate the terminal value of the company using a PE multiple. The industry PE multiple is 13. What is your new estimate of the company’s stock price?
Answer #1

a.Calculation of free cash flow for year 1 = (Sales - Costs) x (1-t) - Net Investment

                                                              = (148.2 million - 82.6 million) x (1-0.39) - 16.6 million

                                                              = 23.416 million i.e. 23,416,000

Free Cash flow from Year 1 to Year 6

Year

Growth Rate

Value

Year 1

$23,416,000

Year 2

14%

$26,694,240

Year 3

12%

$29,897,549

Year 4

10%

$32,887,304

Year 5

8%

$35,518,288

Year 6

6%

$37,649,385

Year 7

6%

$39,908,348

Now we will calculate the value of Maslyn Manufacturing by Present value of Free cash flows

After the year 5 growth rate will stabilize at 6%. So we will calculate terminal value at the end of year 5.

Terminal value at year 5 = Cash flows for year 6 / (Ke - g)

                                       = 37,649,385/(13% - 6%)

                                       = $537,848,357

PV of Cash flows from year 1 to year 5 = 23,416,000/1.13 + 26,694,240/1.132 + 29,897,549/1.133 + 32,887,304/1.134 + 35,518,288/1.135

                                                              = $101,796,433

Total PV of Cash Flows = $537,848,357/1.135 + 101,796,433

                                      = 291,922,540 + 101,796,433

                                     = $393,718,973

Value of each share = Value of Company/No of shares

                                = 393,718,973/7,100,000

                                = $55..45

b. EPS per share for year 5 = Earnings for year 5/No. of shares

                                            = $35,518,288/7,100,000

                                            = 5.0026

Terminal value using PE multiple = EPS x PE multiple = 5.0026 x 13 = 65.03

PV of terminal value per share = 65.03/1.135 = 35.30

PV of cash flows from year 1 to 5 = 101,796,433

PV of cash flows per share = 101,796,433/7,100,000 = $14.34

So, Value per share = 35.30 + 14.34 = $49.64